OpenStudy (anonymous):

I will give medal if answer . 3x + 2y =6

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

not enough info

OpenStudy (anonymous):

Graph 3x + 2y = 6 3x-3x + 2y =-3x+ 6 [solve for y] 2y=-3x+6

OpenStudy (anonymous):

which of the ordered pairs in the form (x,y) is a solution of this equation

OpenStudy (anonymous):

x=2-2y/3